We are using sequential search to search an array of size n. It is known that the item we are looking is definitely present in the array. The probability that the item we are looking for is the last one in the array is 1/3. The probabilities of all other items are equal. What is the average case time complexity(counting the number of comparisons) of the algorithm in this case?
Answer #1

`Hey,

Note: Brother if you have any queries related the answer please do comment. I would be very happy to resolve all your queries.

Since array has length n

So, Expected time is

(1/3)*n+(2/(3*(n-1)))*((n-1)+(n-2)+.....1)

E=n/3+(n*(n-1))/(3*(n-1))=n/3+n/3=2n/3

So, Expected time is

2n/3

Kindly revert for any queries

Thanks.